Two Electric field vectors E1 and E2 are perpendicular to each other; obtain its base
vectors.
1
Expert's answer
2020-05-14T14:37:47-0400

Imagine that these two vectors make up a resulting vector E θ degrees above i-axis, therefore, we can write that


"\vec{E}=|E|\text{cos}\theta \textbf{i}+|E|\text{sin}\theta \textbf{j}."

Now put that


"a=\frac{|E_2|}{|E|\text{cos}\theta},\space b=\frac{|E_1|}{|E|\text{sin}\theta}."

Thus,


"\vec{E}=a|E_2|\textbf{i}+b|E_1|\textbf{j}."

Now the vector E can be written as a linear combination of "\vec{E_1},\vec{E_2}." The basis vector is


"<\vec{E_1},\vec{E_2}>."